2.1 Consider the production function: Y=2X + 16X2- 0.4X3 a) Derive marginal value product (VMP) and average value product (AVP) equations. [04] b) At what levels of X do the TPP reach its maximum? [03] c) At what level of X does stage two (II) begin or stage one (I) end? [04] d) Find and interpret elasticity of production (EP) when X = 1. [02] 2.2 Market structures are referred as the competitive environment in which the buyers and sellers of the product operate. A market consists of all the actual and potential buyers and sellers of a particular product. Discuss thoroughly oligopolistic and perfect competitive market structure and give examples of each form of the market structure. [08] 2.3 Differentiate between product market and labour market (Use diagrams that explain the interactions in the markets). [04


1
Expert's answer
2022-02-21T11:55:21-0500

"Y=2X+16X^{2}-0.4X^{3}"

"VMP=\frac{dy}{dx}=2+32X-1.2X^{2}"

"AVP=\frac{Y}{X}=2+16X-0.4X^{2}"

TPP is maximum when VMP is 0

"2+32X-1.6X^{2}=0"

Solving using quadratic formula

"X=\frac{-b+-√(b^{2}-4ac)}{2a}"

"=\frac{-32+-√(32^{2}-4×-1.2×2)}{2×-1.2}"

"X=26.7"

Stage 2 begins when"VMP=AVP"

"2-32X-1.2X^{2}=2+16X-0.4X^{2}"

"=0-48X-0.8X^{2}"

Solving for X using quadratic formula

"X=\frac{48+-√(48^{2}-4×0.8×0)}{2×-0.8}=60"

"EP=\frac{VMP}{AVP}"

"=\frac{2-32X-1.2X^{2}}{2+16X-0.4X^{2}}=-1.8"

The products demand is inelastic since the EP is less than 1.